OpenStudy (anonymous):

What is the solution set of 2x(x - 1) = 3?

OpenStudy (ash2326):

\[2x(x - 1)=3\] Let's simplify this, I'm gonna open the bracket on the left hand side \[2x^2-2x=3\] so we get \[2x^2-2x-3=0\] Do you know quadratic formula?

OpenStudy (anonymous):